The corroded contacts in a lightbulb socket have 8.3 Ω total resistance.

Part A

How much actual power is dissipated by a 110 W (120 V) lightbulb screwed into this socket?

Express your answer using two significant figures.
Answer #1

given,

resistance = 8.3 ohm

power = 110 W

voltage = 120 V

power = voltage^2 / resistance

110 = 120^2 / resistance

resistance of the bulb = 130.91 ohm

total resistance = resistance of the bulb + resistance of the socket

total resistance = 130.91 + 8.3

total resistance = 139.21 ohm

power = 120^2 / 139.21

actual power is dissipated = 103.441 watt